OpenStudy (anonymous):

How do you figure out the limit of sin^2(5x)/(5x^2)

OpenStudy (freckles):

x approaches?

OpenStudy (anonymous):

as x approaches 0, sorry, forgot that!

OpenStudy (freckles):

\[\\ \lim_{u \rightarrow 0} \frac{\sin(u)}{u}=1 \\\] --- \[\lim_{x \rightarrow 0} \frac{\sin^2(5x)}{5x^2} \\ \lim_{x \rightarrow 0} \frac{\sin(5x) \cdot \sin(5x)}{5x^2} \\ \lim_{x \rightarrow 0} \frac{\sin(5x)}{5x} \cdot \frac{\sin(5x)}{5x} \cdot 5\]

OpenStudy (freckles):

could you continue from here?

OpenStudy (anonymous):

yes, thank you!